If a resistance is considerable what is the relation between time of Ascent and time of descent

Grade:11

1 Answers

Khimraj
5 years ago
let air resistance is a then
time of ascent  ta = \sqrt{2h/(g+a)}
time of descent td = \sqrt{2h/(g-a)}
ta/td = \sqrt{(g-a)/(g+a)}
Hope it clears...............................................................